Skip to content
doubtly

In a town 10 accidents take place in the span of 50 days. Assuming that the number of accidents follows Poisson distribution, find the probability that there will be 3 or more accidents on a day. [ Given that \(e^{-0.2}=0.8187\) ]Solution : Here \(m=…

Q.6 (C) (ii) In a town 10 accidents take place in the span of 50 days. Assuming that the number of accidents follows Poisson distribution, find the probability that there will be 3 or more accidents on a day. [ Given that \(e^{-0.2}=0.8187\) ]

Solution : Here \(m=\square\) and \(X\sim P(m)\) with parameter m. The p.m.f. of X is : \(P(X=x)=\frac{e^{-m}\cdot m^x}{x!},\ x=0,1,2,\ldots\)

\(P(X\ge 3)=1-P(X\lt 3)=1-\left[\square+\square+\square\right]=1-\left[\frac{e^{-0.2}(0.2)^0}{0!}+\frac{e^{-0.2}(0.2)^1}{1!}+\frac{e^{-0.2}(0.2)^2}{2!}\right]\)

\(=1-[0.8187(1+0.2+0.02)]=1-\square=\square\)

1 answer

Best answer

Answer: \(P(X\ge 3)=0.001186\)

\(m=\) average number of accidents per day \(=\frac{10}{50}=0.2\)

\(P(X\ge 3)=1-[P(X=0)+P(X=1)+P(X=2)]\)

\(=1-\left[\frac{e^{-0.2}(0.2)^0}{0!}+\frac{e^{-0.2}(0.2)^1}{1!}+\frac{e^{-0.2}(0.2)^2}{2!}\right]\)

\(=1-0.8187(1+0.2+0.02)=1-0.8187\times 1.22=1-0.998814=0.001186\)

Doubtly AI · 28 Oct 2026 AI-verified · Oct 2026 Report
Was this answer helpful?

Your answer

Log in or create a free account to answer this question.